OpenStudy (anonymous):

To make jungle juice, Lex mixed 2 3/4 cups pineapple juice, 4 1/2 cups orange juice, 2 cups of grape juice and a cup of vodka. How many cups of jungle juice did she make?

OpenStudy (anonymous):

What kind of school asks this?

OpenStudy (anonymous):

10 1/4

OpenStudy (anonymous):

I swear some humans *slaps hand to forhead shakes head.*

OpenStudy (texaschic101):

lol....they added vodka ok...lets start adding first lets just deal with the whole numbers 2 + 4 + 2 + 1 = 9 cups now the fractions.. 3/4 + 1/2 = 3/4 + 2/4 = 5/4 ---- 5/4 = 1 1/4 cups added together, they equal : 10 1/4 cups